OpenStudy (anonymous):

find the F (the anti-derivative?) of f''(x)=3/(x)^(1/2), where at f'(4)=7 and F(4)=20

OpenStudy (bahrom7893):

yay integration! finally something interesting..

OpenStudy (anonymous):

we cant integrate yet, my calc instructor wont allow it. He wants us to find the anti derivatives

OpenStudy (bahrom7893):

f"(x) = 3/x^(1/2) = 3*x^(-1/2), integrate this.

OpenStudy (bahrom7893):

Integral is anti derivative..

OpenStudy (bahrom7893):

Integral of (3x^(-1/2)) = 3x^((-1/2)+1)/((-1/2)+1)

OpenStudy (bahrom7893):

f'(x) = 3x^(1/2)/(1/2) = 6x^(1/2) + C

OpenStudy (bahrom7893):

f'(4)=7 => 6*sqrt(4) + C = 7 6*2 + C = 7 12+c = 7, c = 5

OpenStudy (bahrom7893):

f'(x) = 6x^(1/2) + 5

OpenStudy (bahrom7893):

repeat the process: integrate (f'(x)) => get _________ + C plug in F(4) = 20 and get the original F

OpenStudy (anonymous):

thank you.